# “My life is miserable, have to sign 500 autographs everyday”: Exposing Humblebragging, the Brags in Disguise

### v1

@@ Abstract
Humblebragging is a phenomenon where individuals present self-promotional statements under the guise of modesty or complaints. For example, a statement like, “Ugh, I can’t believe I got promoted to lead the entire team. So stressful!”, subtly highlights an achievement while pretending to be complaining. Detecting humblebragging is important for machines to better understand the nuances of human language, especially in tasks like sentiment analysis and intent recognition. However, this topic has not yet been studied in computational linguistics. For the first time, we introduce the task of automatically detecting humblebragging in text. We formalize the task by proposing a 4-tuple definition of humblebragging and evaluate machine learning, deep learning, and large language models (LLMs) on this task, comparing their performance with humans. We also create and release a dataset called HB24, containing 3,340 humblebrags generated using GPT-4o. Our experiments show that detecting humblebragging is non-trivial, even for humans. Our best model achieves an F1-score of 0.88. This work lays the foundation for further exploration of this nuanced linguistic phenomenon and its integration into broader natural language understanding systems.
@@
Sharath Naganna1, Saprativa Bhattacharjee1, Pushpak Bhattacharyya1, Biplab Banerjee2, 1Computer Science and Engineering, Indian Institute of Technology Bombay, India 2Centre of Studies in Resources Engineering, Indian Institute of Technology Bombay, India {sharathhn, saprativa, pb}@cse.iitb.ac.in, bbanerjee@iitb.ac.in
$\ast$\$$\ast$\$footnotetext: Equal Contributions
@@ 1 Introduction
Humblebragging is a nuanced socio-linguistic phenomenon where individuals subtly boast about their achievements, possessions, or qualities while disguising their self-promotion with expressions of complaint or humility. The term was coined by American comedian Harris Wittels in 2010, who later published a book on the phenomenon titled Humblebrag: The Art of False Modesty Wittels (2012).
To address why people resort to humblebragging, Trivedi and Srinivas (2019) analyzed its rising popularity, particularly on social media, from the dual perspective of need theories (Maslow’s Hierarchy of Needs, Maslow and Lewis 1987; see footnote 2) and the Hubris Hypothesis Hoorens et al. (2012). Humblebragging allows a person to satisfy the dual needs of belonging (level 3) through humility and self-esteem (level 4) through bragging. Moreover, the hubris hypothesis states that addressees prefer implicit self-superiority claims over explicit ones. Straightforward bragging also violates the maxims of modesty Leech (1983) and self-denigration Gu (1990), as noted by Zuo (2023). Humblebragging thus serves as a subtle way to convey positive information about oneself, often through expressions of modesty or self-deprecation. Even though it is commonly observed on social media (Twitter, Reddit, and Instagram), many people also use it in real-life conversations Wittels (2012); Sezer et al. (2018); Trivedi and Srinivas (2019). Examples of humblebrags are provided in Table 1.
@@
Humblebragging is closely related to other forms of figurative language, such as sarcasm Gibbs (1986) and irony Garmendia (2018). All three rely on verbal incongruity, where there is a contrast between what is being said and what is actually meant. While extensive research exists on the computational modeling of sarcasm Bhattacharyya and Joshi (2017); Riloff et al. (2013); Cai et al. (2019) and irony Zeng and Li (2022); Van Hee et al. (2018); Barbieri and Saggion (2014), to the best of our knowledge, this work is the first to explore humblebragging in computational linguistics
@@
    Introducing the task of automatic humblebragging detection from text to the computational linguistics community by proposing a 4-tuple definition of humblebragging for streamlining its processing (section 3).
@@
    Benchmarking various machine learning, deep learning, and state-of-the-art large language model techniques on the task of automatic humblebragging detection from text (section 5 and section 6).
@@
    Releasing a new dataset named HB24 on humblebragging detection, containing 3,340 humblebrags, to enable further research on the task (section 4).
@@ 2 Related Work
Both sarcasm and irony have been extensively studied in computational linguistics over the past two decades. Joshi et al. (2015) demonstrated how incongruences can enhance sarcasm detection, while Joshi et al. (2017a) categorized detection methods, benchmark datasets, and evaluation metrics. Recently, Gole et al. (2023) explored the use of large language models for sarcasm detection. Beyond detection, Joshi et al. (2017b) proposed a hybrid rule-based and statistical approach for identifying sarcasm targets, complemented by transformer-based methods like BERT, as shown by Parameswaran et al. (2021). For irony, Zeng and Li (2022) provided a comprehensive survey on computational approaches. Hernández-Farías et al. (2015) evaluated traditional machine learning models for irony detection using sentiment scores, while Wen et al. (2023) introduced the Retrieval–Detection Method for Verbal Irony (RDVI), leveraging open-domain resources for enriched detection.
@@
Finally, there has been a growing trend towards performing classification tasks by posing them as generation tasks. This approach to text classification is particularly relevant for leveraging decoder-based large language models in classification settings, where text generation mechanisms complement traditional methods. For instance, LLMs have been employed as zero-shot Gretz et al. (2023) and few-shot Mirza et al. (2024) text classifiers. Moreover, Saunshi et al. (2021) provides mathematical insights into modeling classification tasks as text completion tasks
3 Task Formulation
In this section, we formally define humblebragging and outline the associated tasks.
@@ 3.1 Formal Definition of Humblebragging
Similar to the 6-tuple definition of sarcasm proposed by Ivanko and Pexman (2003), we define humblebragging as a 4-tuple to systematically capture its key components and underlying structure:
@@
  • $BT$: Brag Theme – The overarching theme or specific category of the brag embedded within the statement. Categories are listed in Appendix G.
  • $HM$: Humble Mask – The segment of the text that adopts a modest or self-deprecating tone to obscure or mitigate the act of bragging.
  • $MT$: Mask Type – Specifies whether the humble mask adopts a modest tone or a self-deprecating approach.
@@
3.2 Binary Classification for Humblebragging Detection
In the context of humblebragging detection, the task is to classify a given text $x$ as either $C_{\text{HB}}$ (Humblebragging) or $C_{\text{Non-HB}}$ (Non-Humblebragging). The process involves generating text encodings from the input, which are then used for classification. For further details, we refer the reader to Appendix A.
Though decoder models are primarily designed for language generation tasks, their ability to predict the next token in a sequence makes them adaptable to various natural language understanding tasks, including classification. Humblebragging classification can be reformulated as a Yes/No question-answering task, where the model determines whether the input text contains humblebragging. This approach leverages the natural language understanding capabilities of pre-trained language models to classify text.
Framework
The input text $x$ is transformed into a prompt structured as a question. For example:
$<question><x><answer>$
The model is given the $question$ and $x$ as the input prompt and is expected to generate a text completion for the $answer$. The $answer$ is then analyzed to determine whether it contains the required word.
  • If $answer$ contains "Yes", the input is classified as $y=1$ (Humblebragging).
  • If $answer$ contains "No", the input is classified as $y=0$ (Non-Humblebragging).
@@ 4 Dataset
As there were no existing datasets for the task of humblebragging detection, we propose HB24, a well-balanced collection of humblebrag and non-humblebrag texts, comprising both human-written and synthetic samples. Due to the limited availability of quality humblebrags, we leverage the capabilities of large language models to generate human-like examples, augmenting the existing data and enhancing the corpus for training classification models. The synthetic data is used for training, while the trained model’s performance is evaluated on human-written samples. Table 2 presents the dataset statistics.
@@
Min 1 6
Max 70 47
Avg 19.55 17.5
Table 2: Dataset statistics. Min, Max and Avg refers to the minimum, maximum and average number of words respectively.
@@ 4.1 Human-Written Humblebrags
Wittels (2012) presents a curated collection of high-quality humblebrag tweets, categorized into themes such as wealth, first-class travel, workplace achievements, celebrity status, and more. These tweets form the positive class within our test set.
@@ 4.2.1 Zero-Shot Prompts
In the zero-shot generation setup, we used two types of prompts. In the General Prompt, we did not explicitly define humblebragging; instead, we asked the model to generate tweets that subtly mention various achievements. In the Prompt with Themes, we provided a formal definition of humblebragging along with the themes (Appendix G) outlined in Wittels (2012). The prompts are provided in Appendix B.
@@ 4.2.3 Post-Processing and Data Curation
After executing all three prompts, we generated a total of 11,000 synthetic samples containing humblebrags. Each sample was manually reviewed to assess its quality and relevance. From this pool, we filtered tweets from each prompt type, ensuring a balanced selection (see Table 3).
@@ 4.3 Non-Humblebrags
Humblebrags are often confused with sarcasm and irony, as all involve an incongruence between the utterance and its intended meaning. To help the model distinguish these phenomena, we included sarcasm and irony as negative samples, alongside direct brags and straightforward complaints. Direct brags convey explicit self-promotion, while complaints reflect surface emotions often present in humblebrags.
@@ 5.2 Encoder Models
Two transformer-based Vaswani et al. (2017) encoder models, BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), were evaluated on the task of humblebragging classification. The experiments utilized the Adam optimizer, and 5-fold cross-validation was employed for hyperparameter tuning. Details of the hyperparameters and training configurations are provided in Appendix E.
@@ 5.3 Decoder Models
For the task of detecting humblebragging, decoder models were evaluated in three settings: zero-shot classification (Z), zero-shot with the 4-tuple definition of humblebragging (Z+D), and fine-tuning on the humblebragging dataset (F). In Z, models classify statements as humblebrag or not, relying only on the input. In Z+D, models were contextually grounded using the 4-tuple definition introduced in subsection 3.1. The prompts are mentioned in Appendix D.
We used open-source models namely, Qwen2.5-7B-Instruct Yang et al. (2024), Llama 3.1-8B-Instruct Dubey et al. (2024), Gemma 1.1 7B (IT) Team et al. (2024), Mistral-7B-Instruct-v0.3 Zheng et al. (2023a), and Vicuna-7B-v1.5 Zheng et al. (2023b) from Hugging Face’s pre-trained model repository [^3] https://huggingface.co/, along with proprietary models GPT-3.5 and GPT-4o. Model outputs were limited to two tokens, and each model was evaluated five times to record average metrics. GPT-3.5 and GPT-4o predictions were obtained via OpenAI’s API [^4] https://platform.openai.com/docs/overview. Hyperparameter details are provided in Appendix E.
5.4 Human Performance
5.5 Evaluation Metrics
We evaluate model performance using Macro Precision, Macro Recall, Macro F1-Score, and Accuracy, ensuring a comprehensive analysis of the results.
@@ 6.1 Quantitative Analysis
Model Precision Recall F1-Score Accuracy
Baseline 0.25 0.50 0.34 0.51
Human 1 0.89 0.81 0.85 0.86
Human 2 0.86 0.81 0.84 0.84
Human 3 0.82 0.51 0.63 0.70
Average 0.86 0.71 0.77 0.80
Logistic Regression 0.68 0.58 0.53 0.59
SVM 0.72 0.61 0.56 0.62
BERT-Large-Uncased (F) 0.76 0.50 0.61 0.68
RoBERTa-Large (F) 0.91 0.62 0.74 0.78
GPT-4o (Z) 0.78 0.94 0.85 0.84
GPT-4o (Z+D) 0.91 0.85 0.88 0.89
GPT-3.5 (Z) 0.65 0.60 0.57 0.61
GPT-3.5 (Z+D) 0.76 0.75 0.75 0.75
Qwen2.5-7B-Instruct (Z) 0.82 0.35 0.49 0.64
Qwen2.5-7B-Instruct (Z+D) 0.85 0.50 0.63 0.71
Qwen2.5-7B-Instruct (F) 0.85 0.40 0.54 0.67
Mistral-7B-Instruct-v0.3 (Z) 0.55 0.96 0.70 0.60
Mistral-7B-Instruct-v0.3 (Z+D) 0.55 0.96 0.70 0.60
Llama-3.1-8B-Instruct (Z) 0.49 0.99 0.66 0.49
Llama-3.1-8B-Instruct (Z+D) 0.62 0.88 0.72 0.68
Llama-3.1-8B-Instruct (F) 0.87 0.72 0.79 0.81
@@
Gemma-1.1-7b-it (Z+D) 0.53 0.83 0.65 0.56
Gemma-1.1-7b-it (F) 0.71 0.44 0.60 0.71
Vicuna-7b-v1.5 (Z) 0.60 0.28 0.38 0.55
Vicuna-7b-v1.5 (Z+D) 0.62 0.51 0.56 0.61
@@
Overall, the best-performing model in terms of F1-score is GPT-4o (0.88 F1), outperforming even the best human annotator (0.85 F1). We speculate this might be due to the vast linguistic and world knowledge acquired by such large parameter models. Among the three human annotators, one performed significantly worse than the others, indicating that the task is non-trivial and may even be difficult for some individuals. Moreover, two notable observations emerge from the results.
First, across all decoder models, Z+D versions consistently outperform their corresponding Z versions. This demonstrates that our 4-tuple definition aids the models in identifying instances of humblebragging. To further verify that the 4-tuple definition enhances the models’ understanding of humblebragging, we conducted additional experiments: (a) replacing the definition in the system prompt with a completely random gibberish text (Appendix D) and (b) replacing our 4-tuple definition with an alternative definition [^5] We refer to this as the textbook definition (TD). provided by Wittels in his book Wittels (2012).
A specific type of brag that masks the boasting part of a statement in a faux-humble guise. The false humility allows the offender to boast about their “achievements” without any sense of shame or guilt. Humblebrags are usually self-deprecating in nature.
The results presented in Figure 2 clearly demonstrate that our 4-tuple definition significantly improves the models’ performance as humblebrag classifiers.
@@
Second, fine-tuning with our HB24 dataset improved the F1-scores of the majority of the models. Both encoder models and three decoder models (Llama, Gemma, and Qwen) showed significant gains from fine-tuning. Interestingly, fine-tuned RoBERTa outperformed all 7–8 billion parameter decoder models except for Llama (F). This highlights the superior classification capabilities of encoder-only models when a high-quality dataset for fine-tuning is available.
We illustrate in Figure 3 Llama’s progression from an F1-score of 0.66 in the zero-shot setting (Z) to an F1-score of 0.79 after fine-tuning (F), through confusion matrices. Llama (Z) primarily predicted the yes label for almost all samples. With our 4-tuple definition in Llama (Z+D), the model began to identify non-humblebrags, bringing more balance to the confusion matrix. After fine-tuning, Llama (F) became more proficient in identifying non-humblebrags while sacrificing some true positives.
Lastly, we observed anomalous behavior for Mistral and Vicuna, where fine-tuning resulted in a decrease in F1-scores. For Mistral, one possible explanation could be that its zero-shot (Z) performance in humblebragging classification was already much higher than all other models in its category and even exceeded that of the much larger GPT-3.5. In the case of Vicuna, the fine-tuned model generated random texts and emojis. Also we observed very high sensitivity to minimal changes in prompts in Z+D setting. The same prompt used for other models resulted in a drop in F1-score. We had to remove the last sentence from the system prompt (Appendix D) which improved the performance. This behavior of Vicuna demands further investigation.

### v2

@@ Abstract
Humblebragging is a phenomenon in which individuals present self-promotional statements under the guise of modesty or complaints. For example, a statement like, “Ugh, I can’t believe I got promoted to lead the entire team. So stressful!”, subtly highlights an achievement while pretending to be complaining. Detecting humblebragging is important for machines to better understand the nuances of human language, especially in tasks like sentiment analysis and intent recognition. However, this topic has not yet been studied in computational linguistics. For the first time, we introduce the task of automatically detecting humblebragging in text. We formalize the task by proposing a 4-tuple definition of humblebragging and evaluate machine learning, deep learning, and large language models (LLMs) on this task, comparing their performance with humans. We also create and release a dataset called HB-24, containing 3,340 humblebrags generated using GPT-4o. Our experiments show that detecting humblebragging is non-trivial, even for humans. Our best model achieves an F1-score of 0.88. This work lays the foundation for further exploration of this nuanced linguistic phenomenon and its integration into broader natural language understanding systems.
@@
Sharath Naganna*, Saprativa Bhattacharjee*, Biplab Banerjee, Pushpak Bhattacharyya Indian Institute of Technology Bombay, Mumbai, India {sharathhn, saprativa, pb}@cse.iitb.ac.in, bbanerjee@iitb.ac.in
$\ast$\$$\ast$\$footnotetext: Equal contribution.
@@ 1 Introduction
Humblebragging is a nuanced socio-linguistic phenomenon in which individuals subtly boast about their achievements, possessions, or qualities while disguising their self-promotion with expressions of complaint or modesty. The term was coined by American comedian Harris Wittels in 2010, who later published a book on the phenomenon titled Humblebrag: The Art of False Modesty Wittels (2012). A few examples of humblebrags are provided in Table 1.
@@
To address why people resort to humblebragging, Trivedi and Srinivas (2019) analyzed its rising popularity, particularly on social media, from the dual perspectives of need theories (Maslow’s Hierarchy of Needs, Maslow and Lewis 1987; see footnote 2) and the Hubris Hypothesis Hoorens et al. (2012). Humblebragging allows a person to satisfy the dual needs of belonging (level 3) through humility; and self-esteem (level 4) through bragging. Moreover, the hubris hypothesis states that addressees prefer implicit self-superiority claims over explicit ones. Also, straightforward bragging violates the maxims of modesty Leech (1983) and self-denigration Gu (1990), as noted by Zuo (2023). Humblebragging thus serves as a subtle way to convey positive information about oneself, often through expressions of modesty or complaint. Even though it is commonly observed on social media (Twitter, Reddit, and Instagram), many people also use it in real-life conversations Wittels (2012); Sezer et al. (2018); Trivedi and Srinivas (2019). Although humblebragging has been recognized socially and culturally, it remains unexplored through the lens of computational linguistics. Evidence from search trends, social media behavior, and historical usage suggests that humblebragging is a persistent and recognizable phenomenon, even when not explicitly labeled. We provide supporting statistics and examples in Appendix A.
@@
Humblebragging is closely related to other forms of figurative language, such as sarcasm Gibbs (1986) and irony Garmendia (2018), which also rely on verbal incongruity, a contrast between what is said and what is meant. While irony typically contrasts expectation with reality and sarcasm adds a mocking tone, humblebragging uniquely conceals self-promotion within a modest or complaining remark. A detailed comparison of these differences is provided in Appendix B. Although extensive research exists on computational modeling of sarcasm Bhattacharyya and Joshi (2017); Riloff et al. (2013); Cai et al. (2019) and irony Zeng and Li (2022); Van Hee et al. (2018); Barbieri and Saggion (2014), to the best of our knowledge, this work is the first to explore humblebragging in computational linguistics.
@@
    Introduction of the task of automatic humblebragging detection from text to the computational linguistics community by proposing a 4-tuple definition of humblebragging for streamlining its processing (section 3 and section 5).
@@
    Benchmarking of various machine learning, deep learning, and state-of-the-art large language model techniques on the task of automatic humblebragging detection from text (section 6 and section 7).
@@
    Release of a new dataset named HB-24 on humblebragging detection, containing 3,340 humblebrags, to enable further research on the task (section 4).
@@ 2 Related Work
Both sarcasm and irony have been extensively studied in computational linguistics over the past two decades. Joshi et al. (2015) demonstrated how incongruences can enhance sarcasm detection, while Joshi et al. (2017a) categorized detection methods, benchmark datasets, and evaluation metrics. Recently, Gole et al. (2023) explored the use of large language models for sarcasm detection. Beyond detection, Joshi et al. (2017b) proposed a hybrid rule-based and statistical approach for identifying sarcasm targets, which was later complemented by transformer-based methods such as BERT, as demonstrated by Parameswaran et al. (2021). For irony, Zeng and Li (2022) provided a comprehensive survey on computational approaches. Hernández-Farías et al. (2015) evaluated traditional machine learning models for irony detection using sentiment scores, while Wen et al. (2023) introduced the Retrieval–Detection Method for Verbal Irony (RDVI), leveraging open-domain resources for enriched detection.
@@
Finally, there has been a growing trend towards performing classification tasks by posing them as generation tasks. This approach to text classification is particularly relevant for leveraging decoder-based large language models in classification settings, where text generation mechanisms complement traditional methods. For instance, LLMs have been employed as zero-shot Gretz et al. (2023) and few-shot Mirza et al. (2024) text classifiers. Moreover, Saunshi et al. (2021) provides mathematical insights into modeling classification tasks as text completion tasks.
3 Formulation of the Humblebragging Definition
In this section, we define and derive our proposed framework for humblebragging.
@@ 3.1 Formal Definition of Humblebragging
We define humblebragging as a 4-tuple to systematically capture its key components and underlying structure:
@@
  • $BT$: Brag Theme – The overarching theme or specific category of the brag embedded within the statement. Categories are listed in Appendix C.
  • $HM$: Humble Mask – The segment of the text that adopts a modest or complaining tone to obscure or mitigate the act of bragging.
  • $MT$: Mask Type – Specifies whether the humble mask adopts a modest tone or a complaining approach.
@@
3.2 Derivation of 4-tuple Definition
[…]
Unlike sarcasm, where the Speaker (S) directs an utterance toward a Hearer (H) who must infer the intended meaning, humblebragging often lacks a specific hearer. It is frequently self-directed or broadcasted to a broad audience, making S and H unnecessary in this framework. This adaptation preserves the dual-layered meaning from sarcasm while formalizing humblebragging as a strategic blend of self-effacement and self-promotion.
@@ 4 Dataset
As there were no existing datasets for the task of humblebragging detection, we propose HB-24333Available at https://github.com/SharathHN/HB-24, a well-balanced collection of humblebrag and non-humblebrag texts, comprising both human-written and synthetic samples. Due to the limited availability of quality humblebrags, we leverage the capabilities of large language models to generate human-like examples, augmenting the existing data and enhancing the corpus for training classification models. The synthetic data is used for training, while the trained model’s performance is evaluated on human-written samples. In other words, the training set is composed of synthetic humblebrags while the test set consists of human-written humblebrags. The non-humblebrags are all human-written. Table 2 presents the dataset statistics.
@@
Min #words 1 6
Max #words 70 47
Avg #words 19.55 17.5
Table 2: Dataset statistics. Min, Max and Avg refer to minimum, maximum and average respectively.
@@ 4.1 Human-Written Humblebrags
Wittels (2012) presents a curated collection of high-quality humblebrag texts, categorized into themes such as wealth, first-class travel, workplace achievements, celebrity status, and more. These tweets form the positive class within our test set.
@@ 4.2.1 Zero-Shot Prompts
In the zero-shot generation setup, we used two types of prompts. In the General Prompt, we did not explicitly define humblebragging; instead, we asked the model to generate tweets that subtly mention various achievements. In the Prompt with Themes, we provided a formal definition of humblebragging along with the themes (Appendix C) outlined in Wittels (2012). The prompts are provided in Appendix D.
@@ 4.2.3 Post-Processing and Data Curation
After executing all three prompts, we generated a total of 11,000 synthetic samples containing humblebrags. Each sample was manually reviewed to assess its quality and relevance (see Appendix E for more details). From this pool, we filtered tweets from each prompt type, ensuring a balanced selection (see Table 3).
@@ 4.3 Non-Humblebrags
Humblebrags are often confused with sarcasm and irony, as all three involve an incongruence between the utterance and its intended meaning. To help the model distinguish these phenomena, we included sarcasm and irony as negative samples, alongside direct brags and straightforward complaints. Direct brags convey explicit self-promotion, while complaints reflect surface emotions often present in humblebrags. A more detailed discussion on the differences can be found in Appendix B.
@@
5 Methodology
We define two task setups: a standard binary classification task, and a sentence completion formulation where humblebrag detection is cast as a Yes/No question answering problem. This enables effective use of decoder-only language models for classification through natural language prompts.
5.1 Binary Classification for Humblebragging Detection
In the context of humblebragging detection, the task is to classify a given text $x$ as either $C_{\text{HB}}$ (Humblebragging) or $C_{\text{Non-HB}}$ (Non-humblebragging). The process involves generating text encodings from the input, which are then used for classification. For further details, we refer the reader to Appendix F.
Though decoder models are primarily designed for language generation tasks, their ability to predict the next token in a sequence makes them adaptable to various natural language understanding tasks, including classification. Humblebragging classification can be reformulated as a Yes/No question-answering task, where the model determines whether the input text contains humblebragging or not. This approach leverages the natural language understanding capabilities of pre-trained language models to classify text.
Framework
The input text x is transformed into a prompt structured as
<definition><question><x><answer>
The model is given the <question> along with <definition> and <x> as the input prompt and is expected to generate a text completion for the <answer>. The LLM output in <answer> is then analyzed to determine whether it contains the required word.
  • If <answer> contains "Yes", the input is classified as y=1 (Humblebragging).
  • If <answer> contains "No", the input is classified as y = 0 (Non-humblebragging).
We evaluate this framework under two settings: Z (zero-shot) and Z+D (zero-shot + definition). In the Z setting, the definition is considered null, i.e., no external guidance is provided. In the Z+D setting, the 4-tuple definition is prepended to the input prompt to provide the model with additional context or instruction.
Example:
Input to LLM (Z+D setting):
<definition>: HB = <B,BT,HM,MT>
<question>: Is the given text humblebragging or not? Answer in Yes or No only.
<x>: "Can someone tell the awards committee to chill? Running out of shelf space here!"
Output from LLM:
<answer>: Yes
Classification: y=1
Detailed prompts for both settings are provided in Appendix G.
@@ 6 Experimental Setup
We conducted experiments with machine learning classifiers, encoder models, decoder models and compared the performance with those of human annotators.
@@ 6.1 Machine Learning Classifiers
We evaluated logistic regression and support vector machine (SVM) as simple machine learning based baselines to gauge the task difficulty.
@@ 6.2 Encoder Models
Two transformer-based Vaswani et al. (2017) encoder models, BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), were evaluated on the task of humblebragging classification. The experiments utilized the Adam optimizer, and 5-fold cross-validation was employed for hyperparameter tuning.
@@ 6.3 Decoder Models
For humblebrag detection, decoder models were evaluated in zero-shot (Z) and zero-shot with definition (Z+D) settings, and were also fine-tuned using LoRA (F). In the Z setting, models classify statements as humblebrag or not using only the input. In Z+D, they leverage the 4-tuple definition to guide classification.
Further details about the settings, hyperparameters and human annotators can be found in Appendix H.
@@ 7.1 Quantitative Analysis
Model Accuracy Precision Recall F1-Score
Baseline 0.51 0.25 0.50 0.34
Human 1 0.86 0.89 0.81 0.85
Human 2 0.84 0.86 0.81 0.84
Human 3 0.70 0.82 0.51 0.63
Average 0.80 0.86 0.71 0.77
Logistic Regression 0.59 0.68 0.58 0.53
SVM 0.62 0.72 0.61 0.56
BERT-Large-Uncased (F) 0.68 0.76 0.50 0.61
RoBERTa-Large (F) 0.78 0.91 0.62 0.74
GPT-4o (Z) 0.84 0.78 0.94 0.85
GPT-4o (Z+D) 0.89 0.91 0.85 0.88
GPT-3.5 (Z) 0.61 0.65 0.60 0.57
GPT-3.5 (Z+D) 0.75 0.76 0.75 0.75
Qwen2.5-7B-Instruct (Z) 0.64 0.82 0.35 0.49
Qwen2.5-7B-Instruct (Z+D) 0.71 0.85 0.50 0.63
Qwen2.5-7B-Instruct (F) 0.67 0.85 0.40 0.54
Mistral-7B-Instruct-v0.3 (Z) 0.60 0.55 0.96 0.70
Mistral-7B-Instruct-v0.3 (Z+D) 0.60 0.55 0.96 0.70
Llama-3.1-8B-Instruct (Z) 0.49 0.49 0.99 0.66
Llama-3.1-8B-Instruct (Z+D) 0.68 0.62 0.88 0.72
Llama-3.1-8B-Instruct (F) 0.81 0.87 0.72 0.79
@@
Gemma-1.1-7b-it (Z+D) 0.56 0.53 0.83 0.65
Gemma-1.1-7b-it (F) 0.71 0.71 0.44 0.60
Vicuna-7b-v1.5 (Z) 0.55 0.60 0.28 0.38
Vicuna-7b-v1.5 (Z+D) 0.61 0.62 0.51 0.56
@@
Overall, the best-performing model in terms of F1-score is GPT-4o (0.88 F1), surpassing even the best human annotator (0.85 F1). We speculate that this may be due to the extensive linguistic and world knowledge these large-parameter models possess. Among the three human annotators, one performed significantly worse than the others, indicating that the task is non-trivial and can be challenging for some individuals. A detailed discussion on annotation inconsistency is provided in Appendix I. Moreover, two notable observations emerge from the results.
First, across all decoder models, the Z+D versions consistently outperform their Z counterparts, indicating that our 4-tuple definition effectively aids in detecting humblebragging. To further assess the specific contribution of our definition, we conducted two controlled experiments: (a) replacing the definition in the system prompt with random gibberish (Appendix J), and (b) substituting our 4-tuple definition with the ‘textbook definition’ (TD) by Wittels (2012) (Appendix K).
The results, summarized in Figure 2, clearly demonstrate that models prompted with our 4-tuple definition perform better than both the gibberish and textbook alternatives, reaffirming its utility in enhancing humblebrag classification.
@@
Second, fine-tuning with our HB-24 dataset improved the F1-scores of the majority of the models. Both encoder models and three decoder models (Llama, Gemma, and Qwen) showed significant gains from fine-tuning. Interestingly, fine-tuned RoBERTa outperformed all 7–8 billion parameter decoder models except for Llama (F). This highlights the superior classification capabilities of encoder-only models when a high-quality dataset for fine-tuning is available.
We illustrate in Figure 3 Llama’s progression from an F1-score of 0.66 in the zero-shot setting (Z) to an F1-score of 0.79 after fine-tuning (F), through confusion matrices. Llama (Z) primarily predicted the yes label for almost all samples. With our 4-tuple definition in Llama (Z+D), the model began to identify non-humblebrags, bringing more balance to the confusion matrix. After fine-tuning, Llama (F) became more proficient in identifying non-humblebrags while sacrificing some true positives. Confusion matrices of other models can be found in Appendix L.
Lastly, we observed anomalous behavior with Mistral and Vicuna, where fine-tuning led to lower F1-scores. Notably, Mistral’s zero-shot performance already exceeded that of other models in its category, including the larger GPT-3.5 and any further fine tuning is resulting in catastrophic forgetting. In Vicuna’s case, the fine-tuned model produced random texts and emojis and was extremely sensitive to slight prompt changes in the Z+D setting, requiring removal of the final sentence from the system prompt (Appendix G). Additional insights on performance degradation are provided in Appendix M.